OpenStudy (anonymous):

on a right angle what is x if the other two sides are 25 and 25. Pythagorean Theorem.

OpenStudy (johnweldon1993):

Well we can only assume it is  The pythagorean theorem states that the sum of the squares of the 2 sides are equal to the square of the hypotenuse \[\large a^2 + b^2 = c^2\] We know 'a' and 'b' and want to know c \[\large c = \sqrt{a^2 + b^2}\] Just plug in 25 for 'a' and 'b' and solve for 'c'